OpenStudy (anonymous):

Explain, in complete sentences, how you would use the graphing method to solve the following system of equations and what the final solution will look like on the coordinate plane. x + y alt= -6 2x - 3y < 0

OpenStudy (anonymous):

I've not seen y alt before. Can you tell me what it means?

OpenStudy (anonymous):

oh its suppose to mean greaterthan or equal too . sorry :/

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Well you graph each of the lines. DO you know how to do that?

OpenStudy (anonymous):

nope i dont know how to do that .

OpenStudy (anonymous):

To graph x +y =-6 you need to find some points that make it true and then join them up with a line. So if x = 0, what will y = to make it true?

OpenStudy (anonymous):

-6 ?

OpenStudy (anonymous):

yes. So one co-ordinate will be (0,-6) If y = 0, what will x be?

OpenStudy (anonymous):

would it still be -6 ?

OpenStudy (anonymous):

Yes So another point to plot will be (-6,0). Can you draw that on a graph?

OpenStudy (anonymous):

 It should look a bit like this

OpenStudy (anonymous):

With the other equation it might be best to rearrange it in the form y=...

OpenStudy (anonymous):

y = mx+b ?

OpenStudy (anonymous):

yup. Can you do that?

OpenStudy (anonymous):

i can but i dont really remember how to do it .

OpenStudy (anonymous):

2x-3y=0 add 3y to both sides gives you...

OpenStudy (anonymous):

2x=3

OpenStudy (anonymous):

oops - lost the y should be 2x=3y

OpenStudy (anonymous):

ohh , my bad

OpenStudy (anonymous):

Now we want the y without the 3 so what do we do to get rid of it?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

nope sorry 3 y means 3 times y so we have to do the opposite of multiplying to get rid of it. So what should we do?

OpenStudy (anonymous):

wouldnt you multiply the 3y by 2 for it to equal 6y and do 2x multiply by -3 for it to equal -6 and it cancels out ?

OpenStudy (anonymous):

Whatever you do to one side you have to do to the other side. If you multiply by 2 you will get 4x=6y What you need to do is divide by 3 to get rid of it. Then you will have 2/3x=y

OpenStudy (anonymous):

oh okayyy

OpenStudy (anonymous):

 So your two lines graphed should look a bit like this y=2/3x can you graph it?

OpenStudy (anonymous):

Are you still there?

OpenStudy (anonymous):

yeah sorry my computer froze

OpenStudy (anonymous):

So can you graph the two lines now or do you need a bit more help with that bit?

OpenStudy (anonymous):

a need help . i tried doing it but im just confused

OpenStudy (anonymous):

in the first one you plotted the two points, (0,-6) and (-6,0) and joined them up. is that ok?

OpenStudy (anonymous):

yahh thats okay

OpenStudy (anonymous):

Great Then in the second one, we can use the y=mx+c method. y= 2/3x What is m? What is c?

OpenStudy (anonymous):

i dont know what it is

OpenStudy (anonymous):

Where are you, by the way? What country?

OpenStudy (anonymous):

m is the slope. In this equation it is 2/3 c is the intercept. In this equation it is 0. Is that at all familiar?